OpenStudy (anonymous):

Find the intercepts and then use them to graph the equation y= -3-3x

OpenStudy (anonymous):

y= -3-3x y= -3-3(0) y =-3 y= -3-3x 0= -3-3x 3/-3= x x=-1 (0,-3),(-1,0)